# Reborn: Reinforcement-Learned Boundary Segmentation with Iterative Training for Unsupervised ASR

### v1

@@ 3.1.4   Initializing $\pi_{\theta}$ with Behavior Cloning
Before training the segmentation model $\pi_{\theta}$ with RL, we initialize it with behavior cloning (BC). BC uses supervised training to train the segmentation model to predict the boundaries in the speech features. Given speech features $x_{1:T}$, the segmentation model is trained to predict the 0/1 boundary $\hat{b}_{i}$ by using some boundaries $b_{i}$ as the ground truth. The ground truth boundary is the post-processed boundary (Section 3.1.5) obtained using $\pi_{\theta-1}$, the segmentation model learned in the previous iteration. The BC process is a binary classification task that can be trained with cross-entropy loss. In the first Reborn iteration, we use the $k$-means segmentation boundary from wav2vec-U as the BC target.
3.1.5   Boundary Post-Processing
After training the segmentation model until convergence, we use the segmentation model to predict the boundaries of the whole dataset and post-process those boundaries. The post-processing step uses the phoneme prediction model trained in the previous iteration to refine the boundary predictions $\hat{b}_{i}$. Even if some consecutive speech features are split into two segments by the segmentation model, the mean-pooled features of the two segments may yield the same phoneme prediction. In this case, the two segments will be merged into one segment. An illustration of post-processing is shown in Figure 1(c). Post-processing differs from phoneme sequence de-duplication: post-processing modifies the segmentation prediction $\hat{b}_{i}$, while de-duplication modifies the phoneme sequence.
@@ 3.2 Stage 2: Training the Phoneme Prediction Model
The phoneme predictor takes the mean-pooled features of segmental structures $S=[s_{1},s_{2},...,s_{T^{\prime}}]$ and predicts a phoneme sequence $Y=[y_{1},y_{2},...,y_{T^{\prime}}]$. $S$ are obtained by mean-pooling the features $X$ in the same segmental structure based on $[\hat{b}^{{}^{\prime}}_{1},\cdots,\hat{b}^{{}^{\prime}}_{T^{\prime}}]$, where $\hat{b}^{{}^{\prime}}_{i}$ is the post-processed boundaries. We find that it is essential to perform boundary post-processing to stabilize the training in this stage.
After obtaining the post-processed phoneme boundaries, we train the phoneme predictor model using GAN training. The phoneme prediction model is the generator in GAN training. The generator aims to output phoneme predictions that look like real phoneme sequences to fool the discriminator. The discriminator takes in a phoneme sequence, which can be the output of the generator or a phoneme sequence from the unpaired text corpora, and the goal of the discriminator is to distinguish whether the input phoneme sequences are outputs of the generator. The generator and discriminator are updated using the loss in GAN training. In this stage, the parameters of the segmentation model are not updated.
@@ 3.3 Initialization of Reborn
In Stage 1, the segmentation model depends on the phoneme prediction when calculating the rewards. As a result, Reborn cannot work without properly initializing the phoneme prediction model. We use wav2vec-U to train a phoneme prediction model and use it as the initialization for the phoneme prediction model in Reborn. We briefly introduce wav2vec-U in Appendix A for interested readers.
Algorithm 1 Training process of Reborn
1:Initialization: Use wav2vec-U to obtain a phoneme predictor (§ 3.3)
2:while True do
3:     Stage 1: Initialize segmentation model with behavior cloning (§ 3.1.4) and then train using RL (§ 3.1)
4:     Post-process boundaries (§ 3.1.5)
5:     Stage 2: Train phoneme prediction model (§ 3.2)
6:     if no improvement (§ C.2then break
7:     end if
8:end while
@@ 4.1 Dataset
LibriSpeech (Panayotov et al., 2015) is an English speech recognition corpora that contains 960 hours of training data. Following EURO (Gao et al., 2023), we use 100 hours of audio from the train-clean-100 set as the unlabeled speech data. The unlabeled text data does not overlap with the transcription of the unlabeled speech data.
TIMIT (Garofolo, 1993) is another English speech recognition with the human-labeled phoneme boundary. We follow the matched setting, which is more broadly used (Liu et al., 2018; Baevski et al., 2021; Gao et al., 2023), where the speech and text data come from the same set of utterances.
Multilingual LibriSpeech (MLS) (Pratap et al., 2020) is an ASR dataset including German (de), Dutch (nl), French (fr), Spanish (es), Italian (it), and Portuguese (pt). Following (Baevski et al., 2021), we use 100 hours of speech data for each language and use the LM data provided by the dataset as unpaired text.
@@ 4.2 Evaluation Metrics
We use phoneme error rate (PER) and word error rate (WER) to evaluate the ASR performance. If not specified, we use greedy decoding to obtain phoneme-level results. For decoding word-level outputs, we perform WFST decoding (Mohri, 1997; Mohri et al., 2002) using PyKaldi (Can et al., 2018). We leave the details in Appendix C.3.
@@ 4.3 Implementation Details
For English dataset, we use wav2vec 2.0 (Baevski et al., 2020) to extract speech features from speech signals; for MLS, we use XLSR-53 (Conneau et al., 2021) as the feature extractor. The segmentation model is a two-layer one-dimensional CNN model, detailed in Appendix C.1. For LibriSpeech and TIMIT, we train Reborn for two iterations. For MLS, we only train one iteration since we do not find the performance to improve in the second iteration.
@@ 5.1 Main Results
We show the PER and WER of LibriSpeech, TIMIT, and MLS in Table 1, Table 2, and Table 3. We compare Reborn with several prior UASR works: wav2vec-U (Baevski et al., 2021), wav2vec-U 2.0 (Liu et al., 2022), and EURO (Gao et al., 2023). We have the following observations:
Table 1: PER/WER on LibriSpeech using 100 hours speech data. $\dagger$: Our reproduction. (wav2vec-U and wav2vec-U 2.0 only report results of using 960 hours of unlabeled speech).
Approach PER/WER (%) $\downarrow$ dev-clean dev-other test-clean test-other With Oracle Boundary Train from oracle 6.3/12.8 9.7/16.3 6.4/12.7 10.0/16.8 Baseline EURO (HuBERT) 15.2/23.1 20.7/29.3 15.1/22.8 21.1/29.8 ${}^{\dagger}$ 19.3/20.4 22.9/25.6 19.3/21.0 23.2/25.8 wav2vec-U ${}^{\dagger}$ 12.2/17.2 16.3/21.7 12.6/17.7 16.3/22.2 wav2vec-U 2.0 + HMM ST ${}^{\dagger}$ 10.0/15.4 13.1/19.0 10.3/16.0 13.1/19.6 Our Method Reborn 8.3/12.5 11.9/17.6 8.9/13.1 12.5/18.7 Reborn + HMM ST 5.2/9.3 8.5/13.5 5.4/9.6 8.5/13.7
@@
Our experiment on LibriSpeech follows Gao et al. (2023) to use the 100-hour training split of LibriSpeech as the unlabeled speech data. Compared with all the baselines in Table 1, Reborn achieves the lowest PER and WER. wav2vec-U and EURO both use heuristics and hand-crafted rules to obtain the segmental boundaries, while wav2vec-U 2.0 removes the feature segmentation steps. The superior performance of Reborn illustrates the importance of learning the segmental boundaries in UASR. Notably, without using HMM self-training, Reborn already has PER/WER lower than wav2vec-U 2.0 with HMM self-training, which is the prior SoTA method on LibriSpeech.
@@
Self-training uses the phoneme predictor’s prediction as the pseudo-label and trains a new phoneme prediction model using the pseudo-label as the training data. It is commonly used in UASR to boost the performance (Baevski et al., 2021; Liu et al., 2022). In Table 1, we show that Reborn can be integrated with HMM self-training to further lower the PER/WER. We reach a new SoTA on LibriSpeech under the setting of 100-hour speech data, outperforming the prior SoTA in PER and WER by 5% and 6%, respectively. Surprisingly, Reborn with HMM self-training outperforms training wav2vec-U with the oracle boundary. This shows that Reborn can be combined with self-training to improve performance effectively. Due to limited computation resources, we only conduct self-training on LibriSpeech.
Table 2: PER results on TIMIT. The cross-mark () in the greedy-decoding column indicates that an additional LM is used during decoding.
@@
Table 2 shows the PER of TIMIT, and we again find that Reborn achieves the lowest PER compared with all prior UASR methods. This indicates that Reborn not only works on large datasets like LibriSpeech but also works on small datasets like TIMIT, which only contains about 3 hours of audio data for training. Even using greedy decoding only, Reborn outperforms the prior best-performing UASR model (row (e)), which relies on prefix decoding. EURO shows that replacing the feature extractor with WavLM (Chen et al., 2022) (row (e)) outperforms using wav2vec 2.0 (row (d)) as the speech feature extractor. Since Reborn using wav2vec 2.0 already outperforms EURO with WavLM, we leave changing the feature extractor in Reborn as future work.
Table 3: PER/WER on MLS. $\dagger$: Results from Baevski et al. (2021). ***: Our reproduction of wav2vec-U, used as the initialization of the phoneme prediction model in Reborn. $\ddagger$: Results from Liu et al. (2022).
Approach WER (%) $\downarrow$ de nl fr es it pt Avg. ${}^{\dagger}$ 32.5 40.2 39.8 33.3 58.1 59.8 44.0 $*{}^{*}$ 33.9 38.1 37.7 33.1 51.8 59.4 42.3 wav2vec-U ${}^{\ddagger}$ 23.5 35.1 35.7 25.8 46.9 48.5 35.9 Reborn 20.9 26.9 28.2 24.7 39.9 51.5 32.0
Figure 2: Comparing Reborn segmentation (a) with oracle boundaries (c) on an example (”It’s healthier to cook…”) from TIMIT test set. The ground truth phoneme transcription for each segment is given in (b). Silences are denoted with a short dash (-). The red crosses indicate wrong phonemes w.r.t. the ground truth; while the green ticks indicate the corrected ones. The dotted lines illustrate the delayed phenomenon as mentioned in Section 5.2.
@@ 5.2 Boundary Analysis
The core of Reborn is the segmentation model that learns the segmental structure’s boundaries. In this section, we look at how the quality of segmental boundaries affects the performance of phoneme prediction. We compare the boundaries obtained by five methods: (a) the $k$-means segmentation boundary in wav2vec-U, (b) the oracle phoneme boundaries obtained by forced alignment (McAuliffe et al., 2017), which requires paired speech-text data to learn the alignment; (c) the phoneme boundary obtained by the SoTA unsupervised phoneme segmentation method (Strgar & Harwath, 2022); (d) the boundaries learned by the segmentation model of Reborn in the first iteration before post-processing and (e) after post-processing.
Table 4: Boundary evaluation results of different segmentation methods on LibriSpeech test-clean split. The second-last column (Freq.) is the number of segments per second. All the methods shares the same phoneme prediction model trained with the $k$-means segmentation of wav2vec-U.
Boundary method Prec. Rec. F1 Freq. PER (%) (a) Oracle 1.0 1.0 1.0 11.89 13.8 (b) $k$-means 0.64 0.77 0.70 14.27 19.3 (c) Strgar & Harwath (2022) 0.75 0.75 0.75 12.26 23.2 (d) Reborn (no post-processing) 0.57 0.78 0.65 16.29 12.9 (e)  + post-processing 0.73 0.71 0.72 11.63 - (f)  + left-shift 1 frame 0.76 0.74 0.75 11.63 -
@@
First, we focus on the PER in Table 4 when using different segmentation methods. A segmentation method segments the speech feature into segments, and the segment-wise mean-pooled features will be used as input to the phoneme prediction model. In Table 4, all the segmentation methods share the same phoneme prediction model: the phoneme prediction model trained using the $k$-means segmentation of wav2vec-U. We find that simply replacing the $k$-means boundary with the oracle phoneme boundary reduces the PER by 5.5%, showing the imperfect heuristic segmentation is the bottleneck of UASR. Next, even when using the boundary predicted by the SoTA unsupervised phoneme segmentation (Strgar & Harwath, 2022), the PER does not reduce. Conversely, the boundary learned by Reborn achieves the lowest PER. This is because the boundaries learned by Reborn are tailored for the phoneme prediction model.
@@
Next, we discuss the phoneme boundary quality in Table 4. The phoneme boundary quality is evaluated with boundary precision, recall, and F1 with a 20ms tolerance window, following the harsh scheme from Strgar & Harwath (2022). We also report the average number of segments per second (Freq. in Table 4). Interestingly, the boundary F1s of the oracle boundary (row (a)) and Strgar & Harwath (2022) (row (c)) are much higher than the boundary F1 of Reborn (row (d)), but the PER of Reborn is much better. This is because Reborn’s segmentation model predicts more segments than the number of phonemes in the utterance (before post-processing), indicating that it learns some segmental structures smaller than the phonemes. This can be seen from the lower precision (0.57) and higher frequency (16.29) compared to oracle boundaries. However, overly segmenting the speech feature is not problematic because even if the consecutive speech features of the same phonemes are split into two segments by the segmentation model, as long as they have the same phoneme prediction, the duplicated phoneme prediction will be removed in the de-duplication step, and thus not affecting the PER.
While Reborn boundaries initially have low precision, with post-processing, boundary precision can be improved by 0.16. Additionally, after applying boundary post-processing, the number of segmental structures is closer to the oracle. In contrast, we find that applying post-processing does not improve the boundary F1 for all other unsupervised segmentation methods listed in Table 4.
We show an example utterance in Figure 2 with the oracle boundary and Reborn boundary (before post-processing). One can observe that Reborn boundaries produce more segments than the oracle boundaries. We also observe an interesting boundary delay phenomenon: the phoneme boundaries predicted by Reborn sometimes lag behind the oracle phoneme boundaries (dotted lines in Figure 2). This observation motivates us to left-shift the post-processed Reborn boundary by one frame, and we find that the resulting boundary F1 is further improved and matches the boundary F1 of the unsupervised phoneme segmentation SoTA method (Strgar & Harwath, 2022) (row (f) v.s. row (c)).

### v2

@@ 3.1.4 Initializing $\pi_{\theta}$ with Behavior Cloning
Before training the segmentation model $\pi_{\theta}$ with RL, we initialize it with behavior cloning (BC) [44, 45]. BC uses the supervised objective to train the segmentation model to predict the boundaries in speech features. Given speech features $x_{1:T}$, the segmentation model is trained to predict the 0/1 boundary $\hat{b}_{i}$ by using some boundaries $b_{i}$ as the ground truth. The ground truth boundary is the merged boundary (Section 3.1.5) obtained using $\pi_{\theta-1}$, the segmentation model learned in the previous iteration. The BC process is a binary classification task that can be trained with cross-entropy loss. In the first Reborn iteration, we use the $k$-means-based segmentation boundary from wav2vec-U as the BC target.
3.1.5 Boundary Merging
After training the segmentation model until convergence, we use the segmentation model to predict the boundaries of the whole dataset and perform boundary merging. The process uses the phoneme prediction model trained in the previous iteration to refine the boundary predictions $\hat{b}_{i}$. Even if some consecutive speech features are split into two segments by the segmentation model, the mean-pooled features of the two segments may yield the same phoneme prediction. In this case, the two segments will be merged into one segment. An illustration of boundary merging is shown in Figure 1(c). Boundary merging differs from phoneme sequence de-duplication: boundary merging modifies the segmentation prediction $\hat{b}_{i}$, while de-duplication modifies the phoneme sequence.
@@ 3.2 Stage 2: Training the Phoneme Prediction Model
The phoneme predictor takes the mean-pooled features of segmental structures $S=[s_{1},s_{2},\cdots,s_{T^{\prime}}]$ and predicts a phoneme sequence $Y=[y_{1},y_{2},\cdots,y_{T^{\prime}}]$. $S$ are obtained by mean-pooling the features $X$ in the same segmental structure based on $[\hat{b}^{{}^{\prime}}_{1},\cdots,\hat{b}^{{}^{\prime}}_{T}]$, where $\hat{b}^{{}^{\prime}}_{i}$ is the merged boundaries. We find that it is effective to perform boundary merging to stabilize the training in this stage.
After obtaining the merged phoneme boundaries, we train the phoneme predictor model using GAN training. The phoneme prediction model is the generator in GAN training. The generator aims to output phoneme predictions that look like real phoneme sequences to fool the discriminator. The discriminator takes in a phoneme sequence, which can be the output of the generator or a phoneme sequence from the unpaired text corpora, and the goal of the discriminator is to distinguish whether the input phoneme sequences are outputs of the generator. The generator and discriminator are updated using the loss in GAN training. In this stage, the parameters of the segmentation model are not updated, and the generator is initialized from the previous iteration. We discuss the effect of applying boundary merging and parameter initialization in Appendix C.3
@@ 3.3 Initialization of Reborn
In Stage 1, the segmentation model depends on the phoneme prediction when calculating the rewards. As a result, Reborn cannot work without properly initializing the phoneme prediction model. We use wav2vec-U to train a phoneme prediction model and use it as the initialization for the phoneme prediction model in Reborn. We briefly introduce wav2vec-U in Appendix D.
@@ 4.1 Dataset
LibriSpeech [42] is an English speech recognition corpora that contains 960 hours of training data. Following EURO [17], we use 100 hours of audio from the train-clean-100 set as the unlabeled speech data. The unlabeled text data is derived from the remaining 860 hours, which does not overlap with the transcription of the unlabeled speech data.
TIMIT [18] is another English speech recognition with the human-labeled phone boundary. We follow the matched setting, which is more broadly used [31, 4, 17], where the speech and text data come from the same set of utterances.
Multilingual LibriSpeech (MLS) [47] is an ASR dataset including German (de), Dutch (nl), French (fr), Spanish (es), Italian (it), and Portuguese (pt). Following [4], we randomly sample and use 100 hours of speech data for each language and use the LM data provided by the dataset as unpaired text.
@@ 4.2 Evaluation Metrics
We use phoneme error rate (PER) and word error rate (WER) to evaluate the ASR performance. If not specified, we use greedy decoding to obtain phoneme-level results. For decoding word-level outputs, we perform WFST decoding [36, 37] using PyKaldi [8]. We leave the details in Appendix E.7.
@@ 4.3 Implementation Details
For the English datasets, we use wav2vec 2.0 [3] [^1] https://dl.fbaipublicfiles.com/fairseq/wav2vec/wav2vec_vox_new.pt to extract speech features from speech signals; for MLS, we use XLSR-53 [13] [^2] https://dl.fbaipublicfiles.com/fairseq/wav2vec/xlsr_53_56k.pt as the feature extractor. Both of them can be found in fairseq [41]. The 4-gram phoneme LM is derived using KenLM [21]. We describe more details about the data preparation and LM formulation in Appendix E.1, which basically follows wav2vec-U [4]. The segmentation model is a two-layer one-dimensional CNN model, detailed in Appendix E.3. For LibriSpeech and TIMIT, we train Reborn for two iterations. For MLS, we only train one iteration since we do not find the performance to improve in the second iteration.
@@ 5.1 Main Results
We show the PER and WER of LibriSpeech, TIMIT, and MLS in Table 1, Table 2, and Table 3. We compare Reborn with several prior UASR works: wav2vec-U [4], wav2vec-U 2.0 [28], and EURO [17]. We have the following observations:
Table 1: PER/WER on LibriSpeech using 100 hours speech data. $\dagger$: Our reproduction. (wav2vec-U and wav2vec-U 2.0 only report results of using 960 hours of unlabeled speech). HMM ST indicates HMM self-training (Appendix E.2).
Approach PER/WER (%) $\downarrow$ dev-clean dev-other test-clean test-other With Oracle Boundary Train from oracle 6.3/12.8 9.7/16.3 6.4/12.7 10.0/16.8 Baseline EURO (HuBERT) 15.2/23.1 20.7/29.3 15.1/22.8 21.1/29.8 wav2vec-U 19.3/20.4 22.9/25.6 19.3/21.0 23.2/25.8 wav2vec-U 2.0 12.2/17.2 16.3/21.7 12.6/17.7 16.3/22.2 wav2vec-U 2.0 + HMM ST 10.0/15.4 13.1/19.0 10.3/16.0 13.1/19.6 Our Method Reborn 8.3/12.5 11.9/17.6 8.9/13.1 12.5/18.7 Reborn + HMM ST 5.2/9.3 8.5/13.5 5.4/9.6 8.5/13.7
Table 2: PER results on TIMIT. The cross-mark () in the greedy-decoding column indicates that an additional LM (4-gram) is used during decoding. Reborn reaches the best performance with no LM used for decoding, showing that Reborn can benefit from the external LM via RL.
@@
Our experiment on LibriSpeech follows Gao et al. [17] to use the 100-hour training split of LibriSpeech as the unlabeled speech data. Compared with all the baselines in Table 1, Reborn achieves the lowest PER and WER. wav2vec-U and EURO both use hand-crafted rules to obtain the segmental boundaries, while wav2vec-U 2.0 removes the feature segmentation steps. The superior performance of Reborn illustrates the importance of learning the segmental boundaries tailored for the phoneme prediction model. Notably, without using HMM self-training, Reborn already has PER/WER lower than wav2vec-U 2.0 with HMM self-training, which is the prior SoTA method on LibriSpeech. In Appendix C.5, we further apply the Reborn pipeline with speech foundation models other than wav2vec, and we find that Reborn yields notable performance improvement when using HuBERT [22] or WavLM [11] as the speech foundation model, illustrating the generalizability of the pipeline.
@@
Self-training uses the phoneme predictor’s prediction as the pseudo-label and trains a new phoneme prediction model using the pseudo-label as the training data. It is commonly used in UASR to boost the performance [4, 28]. In Table 1, we show that Reborn can be integrated with Hidden Markov Models (HMM) self-training (Appendix E.2) to further lower the PER/WER. We reach a new SoTA on LibriSpeech under the setting of 100-hour speech data, outperforming the prior SoTA in PER and WER by 5% and 6%, respectively. Surprisingly, Reborn with HMM self-training outperforms training wav2vec-U with the oracle boundary. This shows that Reborn can be combined with self-training to improve performance effectively. Due to limited computation resources, we only conduct self-training on LibriSpeech.
Table 3: PER/WER on MLS. $\dagger$: Results from Baevski et al. [4]. *: Our reproduction of wav2vec-U, used as the initialization of the phoneme prediction model in Reborn. $\ddagger$: Results from Liu et al. [28].
Approach WER (%) $\downarrow$ de nl fr es it pt Avg. wav2vec-U 32.5 40.2 39.8 33.3 58.1 59.8 44.0 wav2vec-U 33.9 38.1 37.7 33.1 51.8 59.4 42.3 wav2vec-U 2.0 23.5 35.1 35.7 25.8 46.9 48.5 35.9 Reborn 20.9 26.9 28.2 24.7 39.9 51.5 32.0
Table 4: Boundary evaluation results of different segmentation methods on LibriSpeech test-clean split. The second-last column (Freq.) is the number of segments per second. All the methods share the same phoneme prediction model trained with the $k$-means-based segmentation of wav2vec-U.
Boundary method Prec. Rec. F1 Freq. PER (%) (a) Oracle 1.0 1.0 1.0 11.89 13.8 (b) $k$-means-based 0.64 0.77 0.70 14.27 19.3 (c) Strgar and Harwath [53] 0.75 0.75 0.75 12.26 23.2 (d) Reborn 0.57 0.78 0.65 16.29 12.9
@@
Table 2 shows the PER of TIMIT, and we again find that Reborn achieves the lowest PER compared with all prior UASR methods. This indicates that Reborn not only works on large datasets like LibriSpeech but also works on small datasets like TIMIT, which only contains about 3 hours of audio data for training. Even using greedy decoding only, Reborn outperforms the prior best-performing UASR model (row (e)), which relies on prefix decoding. EURO shows that replacing the feature extractor with WavLM [11] (row (e)) outperforms using wav2vec 2.0 (row (d)) as the speech feature extractor. Since Reborn using wav2vec 2.0 already outperforms EURO with WavLM, we leave changing the feature extractor in Reborn on TIMIT as future work.
@@ 5.2 Boundary Analysis
The core of Reborn is the segmentation model that learns the segmental structure’s boundaries. In this section, we look at how the quality of segmental boundaries affects the performance of phoneme prediction. We compare the boundaries obtained by five methods: (a) the oracle phoneme boundaries obtained by forced alignment [34], which requires paired speech-text data to learn the alignment; (b) the $k$-means-based segmentation boundary in wav2vec-U; (c) the phoneme boundary obtained by the SoTA unsupervised phoneme segmentation method [53] on LibriSpeech; (d) the boundaries learned by the segmentation model of Reborn in the first iteration before boundary merging.
@@
First, we focus on the PER in Table 4 when using different segmentation methods. The PER is obtained by taking the four different segmented features to the identical phoneme prediction model: the phoneme prediction model trained using the $k$-means-based segmentation of wav2vec-U. We find that simply replacing the $k$-means-based boundary with the oracle phoneme boundary reduces the PER by 5.5%, showing the imperfect hand-crafted $k$-means-based segmentation is the bottleneck of UASR. Next, even when using the boundary predicted by the SoTA unsupervised phoneme segmentation [53], the PER does not reduce. Conversely, the boundary learned by Reborn achieves the lowest PER. This is because the boundary learned by Reborn is directly tailored for the phoneme prediction model.
@@
Next, we discuss the phoneme boundary quality in Table 4. The phoneme boundary quality is evaluated with boundary precision, recall, and F1 with a 20ms tolerance window, following the harsh scheme from Strgar and Harwath [53]. We leave the discussion about the different schemes for boundary evaluation in Appendix G. We also report the average number of segments per second (Freq. in Table 4). Interestingly, the boundary F1s of the oracle boundary (row (a)) and Strgar and Harwath [53] (row (c)) are much higher than the boundary F1 of Reborn (row (d)), but the PER of Reborn is much better. This is because Reborn’s segmentation model predicts more segments than the number of phonemes in the utterance (before boundary merging), indicating that it learns some segmental structures smaller than the phonemes. This can be seen from the lower precision (0.57) and higher frequency (16.29) compared to oracle boundaries. However, segmenting the speech feature into smaller units is not problematic because even if the consecutive speech features of the same phonemes are split into two segments by the segmentation model, as long as they have the same phoneme prediction, the duplicated phoneme prediction will be removed in the de-duplication step, and thus not affect the PER.
Table 5: We compare different reward functions and the effect of BC initialization with first iteration results on LibriSpeech test-clean. Row (c) is Reborn.
Ablation PER (%) LM PPL (Stage 1) Stage 1 Stage 2 (a). $R_{\text{ppl}}$ 14.7 12.9 10.0 (b). $R_{\text{ppl}}+R_{\text{edit}}$ 13.8 12.1 10.8 (c). $R_{\text{ppl}}+R_{\text{edit}}+R_{\text{len}}$ 12.9 11.9 11.2 (d). Reborn w/o BC 14.9 14.2 13.8